# Non-classical nucleation pathways in liquid condensation revealed by simulation and theory

### v4

@@ Abstract
Prompted by the confirmations of its quality, our model can be used to predict the work of formation, see Fig. 3. Tolman correction, that is almost imperceptible on the nucleation barrier, is not shown. At low supersaturation, the three models agree because critical cluster density matches closely the equilibrium liquid density. In the intermediate supersaturation region, both our model and SGA predict much smaller nucleation barriers. It is counter-intuitive at first since the driving force for nucleating a dilute cluster is smaller as compared to a cluster with the equilibrium density, but it is balanced by a much lower surface energy cost. Finally, when approaching the spinodal limit, the two-variable nucleation theory recovers results consistent with diffuse-interface models: the density of critical cluster approaches the initial phase density, its size diverges, and the associated work of formation vanishes. This behavior provides a smooth transition between nucleation and spinodal decomposition, where the energy barrier is zero, as established in diffuse-interface models [Cahn-Hilliard-1959] and recovered in previous nucleation models with two parameters [Schmelzer_2000, Schmelzer_2006, Schmelzer_2003, Schmelzer_2011, Baidakov_2000, Philippe_2011_Phil_mag, Philippe_2024, Bonvalet_Phil_mag_2014, Ghosh-2011, Reiss_1976, Schmelzer_2007]. The surface energy contribution also vanishes at the spinodal limit in our two-variable theory since the cluster density approaches that of the initial phase. CNT, however, fails to capture such effects, highlighting the improved physical accuracy of the two-variable model.
@@
Figure 4: Contour plot of the work of formation overlaid with growth trajectories of critical clusters from simulations at $T=0.8\epsilon/k_{\mathrm{B}}$ and $\rho_{0}=0.03\sigma^{-3}$. The work of formation is computed using our two-variable nucleation theory (without Tolman correction), with the saddle point marked by a red hexagram. The red dash line and blue dotted line indicate the predicted nucleation flux direction and steepest decent direction at the saddle point, respectively. Growth trajectories from four independent simulations are represented by colored crosses, with corresponding mean values shown as dots. Error bars denote standard deviations. $(R,\rho)$ coordinates from theory and simulation are normalized by their corresponding critical values.
@@
Finally, we investigate the nucleation flux direction. In CNT, both kinetics and thermodynamics favor cluster growth at equilibrium density. However, in our two-variable theory the nucleation flux is found to deviate from the energy steepest descent direction when supersaturation increases [See Fig. 4]. Four growth trajectories obtained from brute-force simulations are also superimposed to the corresponding free-energy surface at $\rho_{0}=0.03\sigma^{-3}$. These simulations are initialized with the critical cluster shown in Fig. 1(a), with therefore an equal probability ($\sim 0.5$) of growing or shrinking but we keep only growth trajectories for the sake of clarity. From Fig. 4, we observe that contrary to the CNT predictions, the nucleation pathway is not an horizontal line but involves concomitant growth and densification of the cluster. More importantly, the numerical results align more closely with our model than with the steepest descent predictions, providing further confirmation of the quality of the two-variable nucleation theory.
@@
Conclusions - We revisited the longstanding yet still fundamental problem of liquid condensation and uncovered a non-classical nucleation pathway. Using innovative rare-event simulation strategies – an original implementation of the seeding approach together with steering MD combined with aimless shooting – we were able to probe nucleation events with unprecedented precision and collect structures situated at the saddle point of the free energy barrier across a wide range of supersaturation. Our simulation results show that the density of the critical clusters differs markedly from that of the equilibrium liquid and decreases as the vapor density increases. Building on this key insight, we expanded the capillary model by explicitly incorporating the cluster density as a supplementary driving parameter. Without any fitted parameters, this self-consistent two-variable nucleation theory successfully captures the concomitant growth and densification of nearly critical cluster, as observed in our simulations, and also achieves quantitative agreement for nucleation rates and critical cluster densities. Taken together, these results demonstrate that cluster density is an essential collective variable governing nucleation, and that omitting it is a principal reason for the longstanding failures of CNT. By establishing a multidimensional picture of nucleation, our work provides a broadly applicable framework for refining nucleation theories.
@@ Acknowledgements.
Acknowledgments - This work was supported by the ANR TITANS project. Computational resources have been provided by the DSI at Université de Lille, Jean-Zay and TGCC.

### v5

@@ Abstract
Prompted by the confirmations of its quality, our model can be used to predict the work of formation; see Fig. 3. The Tolman correction, which is almost imperceptible on the nucleation barrier, is not shown. At low supersaturation, the three models agree because critical cluster density matches closely the equilibrium liquid density. In the intermediate supersaturation region, both our model and the SGA model predict much smaller nucleation barriers. It is counterintuitive at first since the driving force for nucleating a dilute cluster is smaller as compared to a cluster with the equilibrium density, but it is balanced by a much lower surface energy cost. Finally, when approaching the spinodal limit, the two-variable nucleation theory recovers results consistent with diffuse-interface models: the density of the critical cluster approaches the initial phase density, its size diverges, and the associated work of formation vanishes. This behavior provides a smooth transition between nucleation and spinodal decomposition, where the energy barrier is zero, as established in diffuse-interface models [Cahn-Hilliard-1959] and recovered in previous nucleation models with two parameters [Schmelzer_2000, Schmelzer_2006, Schmelzer_2003, Schmelzer_2011, Baidakov_2000, Philippe_2011_Phil_mag, Philippe_2024, Bonvalet_Phil_mag_2014, Ghosh-2011, Reiss_1976, Schmelzer_2007]. The surface energy contribution also vanishes at the spinodal limit in our two-variable theory since the cluster density approaches that of the initial phase. CNT, however, fails to capture such effects, highlighting the improved physical accuracy of the two-variable model.
@@
Figure 4: Contour plot of the work of formation overlaid with growth trajectories of critical clusters from simulations at $T=0.8\epsilon/k_{\mathrm{B}}$ and $\rho_{0}=0.03\sigma^{-3}$. The work of formation is computed using our two-variable nucleation theory (without Tolman correction), with the saddle point marked by a red hexagram. The red dashed line and blue dotted line indicate the predicted nucleation flux direction and steepest decent direction at the saddle point, respectively. Growth trajectories from four independent simulations are represented by colored crosses, with corresponding mean values shown as dots. Error bars denote standard deviations. $(R,\rho)$ coordinates from theory and simulation are normalized by their corresponding critical values.
@@
Finally, we investigate the nucleation flux direction. In CNT, both kinetics and thermodynamics favor cluster growth at equilibrium density. However, in our two-variable theory the nucleation flux is found to deviate from the energy steepest descent direction when supersaturation increases (see Fig. 4). Four growth trajectories obtained from brute-force simulations are also superimposed to the corresponding free energy surface at $\rho_{0}=0.03\sigma^{-3}$. These simulations are initialized with the critical cluster shown in Fig. 1(a), with therefore an equal probability ($\sim 0.5$) of growing or shrinking but we keep only growth trajectories for the sake of clarity. From Fig. 4, we observe that contrary to the CNT predictions, the nucleation pathway is not an horizontal line but involves concomitant growth and densification of the cluster. More importantly, the numerical results align more closely with our model’s prediction than with the energy steepest descent direction, providing further confirmation of the quality of the two-variable nucleation theory.
@@
Conclusions — We revisited the longstanding yet still fundamental problem of liquid condensation and uncovered a nonclassical nucleation pathway. Using innovative rare-event simulation strategies—an original implementation of the seeding approach together with steering MD combined with aimless shooting—we were able to probe nucleation events with unprecedented precision and collect structures situated at the saddle point of the free energy barrier across a wide range of supersaturation. Our simulation results show that the density of the critical clusters differs markedly from that of the equilibrium liquid and decreases as the vapor density increases. Building on this key insight, we expanded the capillary model by explicitly incorporating the cluster density as a supplementary driving parameter. Without any fitted parameters, this self-consistent two-variable nucleation theory successfully captures the concomitant growth and densification of nearly critical clusters, as observed in our simulations, and also achieves quantitative agreement for nucleation rates and critical cluster densities. Taken together, these results demonstrate that cluster density is an essential collective variable governing nucleation, and that omitting it is a principal reason for the longstanding failures of CNT. By establishing a multidimensional picture of nucleation, our work provides a broadly applicable framework for refining nucleation theories.
@@ Acknowledgements.
Acknowledgments — This work was supported by the ANR TITANS project. Computational resources have been provided by the DSI at Université de Lille, Jean-Zay and TGCC.